### v2

@@ 7 Discussion
The mixed additive relational event model allowed us to formulate a close to generative model of the dynamics driving alien species invasions, by incorporating time-varying and random effects in the model formulation. Dealing with drivers whose effects change over time is indeed quite typical in ecology, especially if the observational time is fairly long as in this context where it covers more than one century. The inclusion of such effects allowed us to understand that the impact of one of the most relevant drivers of alien species invasions, namely trade, is actually decreasing. Furthermore, random effects enable us to account for the heterogeneity of the actors involved in the current and previously occurred events by incorporating network activities. We have included monodic and dyadic random effects in our approach. Particularly, the alien species relational event graph is bimodal, precluding the appearance of triadic network dynamics.
Our estimation method relies on the sampled partial likelihood, which considers in the risk set at each time only one event and non-event. This generates a massive computational advantage, with only minimal information loss. In particular, the computational effort in a Cox Proportional Hazard Model dealing with splines and relying on the partial likelihood scales with $n\times n_{\mathcal{S}}\times n_{\mathcal{C}}\times D$, where $D$ is the dimension of the model matrix of the fitted GAM, including one term for the fixed effect covariates and $q$ terms for the time-varying effect covariates. Instead, case-control partial likelihood via GAMs scales with $n\times D$. We remark that, due to implementation constraints, we were not able to fit a model with $D>n$. For this reason, in our study, when considering the species co-invasion network structure, we did not include all the possible levels that this variable assume in observed events and non-events but we codified dedicated categories. Nevertheless, as a future direction for our efficient technique, we propose sampling more than one non-event for each observed instance. This additional non-event sampling procedure does not have to change the formalization of the logistic regression technique. Instead of only once, the observed occurrence is repeated $m-1$ times, each record with a different sampled non-event. According to this expansion of the method, we are able to increase the number of listed events $n$. At the same time, in presence of random factors with potential large number of levels, we expect to observe more of these levels (leading to an increase of $D$, as well). Concerning the goodness of fit, we tested the proper fitting of the covariates included in the model formulation, both singularly and collectively. Although the test concerning distance led to a rejection, overall the model was judged to be adequate.
@@
Modelling first records meant that we only considered a counting process with zeroes and ones. Novel developments including citizen science initiatives are generating a richer, be it more noisy and more complex, picture of species dispersions. Including a covariate accounting for the number of instances a particular species was detected each year in each country could potentially improve the inference of the drivers of the dynamics of interest (Bonnamour et al., 2021).
[…]
trade $tr_{sr}(t)$ Logarithm of the sum of annual trade flows (\$currency-dollar\\$\$) between region $r$ and other countries that have been invaded by species $s$ before time $t$. This variable shows several missings: our imputation method is described in Section 2.2. Time-varying covariate with time-varying effect (Barbieri et al., 2009)
[…]
Vascular plants 60448 12508 3921 120
Insects and Plants 61546 13094 4035 188
[…]
. Covariate p-value Dimension $q$ distance 0.0002 10 trade 0.580 10 colonial ties 0.404 1 climatic dissimilarity 0.480 1 Omnibus test 0.468
@@ 8 Competing interests
9 Acknowledgments and Funding

### v3

@@ 7 Discussion
The mixed additive relational event model is a rich explanatory and quasi-generative model of the dynamics driving alien species invasions. It is able to integrate and account for various ecological characteristics and socio-economical features. Although not directly comparable, our findings align with various, more qualitative studies in the literature. For instance, Chapman et al. (2017) emphasize the role of trade networks in explaining the large-scale distribution of non-native species, particularly when weighted by proximity and climatic similarity. Similarly, Mwebaze et al. (2023) highlight the interaction between trade and climatic conditions, arguing that trade between countries with similar climates poses a greater invasion risk. These results align with our findings, which identify a positive effect of international trade, measured in a way that incorporates its underlying endogenous network structure. The negative effect of distance suggests that species tend to diffuse over short distances, likely due to a higher probability of survival compared to invasions over larger distances. Moreover, the long-distance dispersal of invasive species is increasingly driven by human-mediated dispersal on a global scale (Zhang et al., 2023). Therefore, the negative impact of distance may be partly explained by the implementation of legal instruments designed to mitigate invasion risks (Turbelin et al., 2017). Greater distances are also often associated with larger climatic differences, making it more difficult for species to establish in distant regions (Morlon et al., 2008; Buckley and Jetz, 2008). This is further supported by the negative effect of climatic dissimilarity, which indicates that species tend to establish in regions with similar climatic conditions, where survival and adaptation are more favorable (Seebens et al., 2015; Juozaitienė et al., 2023). This hypothesis is also reinforced by Liu et al. (2020), who examined ecological niche conservation — a widely accepted concept, albeit with some variability across taxa.
When analyzing first invasion records, we can track the potential occurrence of events at each time point using the risk set. The definition of the risk set at the beginning of the observational period is based on knowledge of species’ native range (Juozaitienė et al., 2023). Elements in the native range cannot be invasion events, by definition. The native range is also used, implicitly, compute several endogenous covariates that depend on knowledge of the regions where a species has already been previously recorded. These covariates, which are central to relational event modeling (Bianchi et al., 2024), incorporate dynamic drivers that go beyond aggregated statistics, usually considered in this framework, such as the distance to the nearest invaded region. As highlighted by Chapman et al. (2017), accounting for a species’ source region, rather than just its destination, is crucial.
Additionally, we have considered endogenous statistics that do not only rely on the source region, but also on other regions where the species might have been present before. For this reason, we argue that our method is able to address the so-called bridgehead effect (Bonnamour et al., 2023). This is a region, where a species has been able to establish, and that serves as a source for new invasive species (Bertelsmeier et al., 2018). Our model is able to incorporate explanatory variables that are computed not only at a species-region level, but also at a temporal level. Furthermore, their effects are also allowed to vary over time. Accounting for drivers whose influence changes over time is particularly relevant in ecology, especially when the observational period spans more than a century, as in our case. As Bonnamour et al. (2021) highlight, the context of globalization has evolved significantly over this period, experiencing two major waves of globalization, leading to substantial changes in transportation systems and, consequently, in trade openness, both of which play a critical role in non-native species introductions. By modeling the effects as varying over time, we confirmed the findings of Juozaitienė et al. (2023), which identified a decreasing impact of trade over time on the hazard of alien species first records. This trend has been attributed, following Luppold (1988) and Mayer et al. (2003), to trade shifts, including changes in commodity structures and the nature of traded goods, leading to a reduction in products that facilitate the spread of alien species.
Our model formulation integrates random effects to account for the heterogeneity of actors involved in both past and present events, as well as to model species co-invasions. In our approach, we include both monadic and dyadic random effects. Through this analysis, we identified Frankliniella occidentalis as the most invasive insect, a finding supported by the literature (Juozaitienė et al., 2023). This is also recognized as a key pest from an economical point of view (Yang et al., 2015). Furthermore, species invasion random effects explicitly enables to account for co-invasion tendencies. This is particularly relevant because the last-invaded species and the newly invading species in a given region may belong to different taxonomic groups. With a similar goal, Bonnamour et al. (2023) examine whether non-native plants facilitate alien insect invasions by providing suitable habitats and resources. They assess this by testing the predictive power of plant flows on non-native insect flows, aggregated over time. Despite using a different methodological approach, our findings align with theirs, showing that the strongest species invasion random effects involve a plant as the last-invaded species and an insect as the newly invading species. Furthermore, from a covariate perspective, the vast majority of last species–species interactions correspond to plant–insect relationships, such as between Phenacoccus manihoti and Chromolaena odorata.
Our estimation method relies on the sampled partial likelihood, which at each event time considers a sampled risk set with only one event and one non-event. This reduces computational complexity, with only minimal information loss. In particular, the computational effort in a Cox Proportional Hazard Model dealing with splines and relying on the partial likelihood scales as $O(n_{\mathcal{S}}n_{\mathcal{C}}nD^{3})$, where $D$ is the dimension of the model matrix of the fitted GAM. Instead, the case-control partial likelihood via GAMs scales as $O(nD^{3})$. The gamm function in the mgcv package is able to add random effects at the cost of 1 degree of freedom, but the disadvantage of this function is its computational cost. Instead, we use the 0-dimensional spline implementation in the gam function in the same package, which is much faster, but which requires by default that $n>D$. Due to this implementation constraint, we did not include all possible levels of the species co-invasion network. Alternatively, it is possible to sample more than one non-event for each observed instance to increase $n$ and to avoid this artificial constraint.
@@
Modelling the high-quality, but binary first records data meant that we only considered a binary counting process. Novel developments, including citizen science initiatives, are nowadays generating a richer, albeit more noisy and more complex, picture of species dispersion. Modelling the number of instances that a particular species was detected each year in each region could potentially improve the picture of the drivers of the dynamics of species invasions (Bonnamour et al., 2021).
@@ 8 Competing interests
9 Author contributions statement
10 Acknowledgments
@@
Data Availability and Reproducibility
The data supporting this study are publicly available at https://zenodo.org/record/4632335. The code required to reproduce the analyses and results is accessible on GitHub at https://github.com/martinaboschi/alienspecies.git.